# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ I.2 Near-Constant Weight Design
To get started, we note that (356) remains valid when $\alpha=0$, which yields that with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$, we have
$n_{2}=\big{(}e^{-\nu(1-\frac{\ell}{k})}+o(1)-\delta^{\prime}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime}|<\delta.$ (374)
@@
Next, we note that (358) remains valid, which together with (374) yields the following with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$:
$M=\big{(}e^{-\nu(1-\frac{\ell}{k})}-e^{-\nu}+o(1)+\delta^{(3)}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{(3)}|<2\delta.$ (375)
@@
Moreover, via the randomness of noise, Hoeffding’s inequality still gives (360), in which we set $t=n_{2}\delta$, substitute (374) and (375), and then perform some simple algebra (as in (361)–(363)) to obtain that
$V=ne^{-\nu(1-\frac{\ell}{k})}\big{(}\rho\star e^{-\frac{\nu\ell}{k}}+o(1)+c_{1,\rho,\nu}\delta\big{)}$ (376)
@@
holds with probability at least $1-2\exp(-e^{-\nu}\delta^{2}n)$, where $c_{1,\rho,\nu}$ is a quantity bounded by a constant only depending on $(\rho,\nu)$. By a union bound, the events (374), (375) and (376) collectively hold with probability at least $1-\exp(-C_{\rho,\nu}\delta^{2}n)$ for some $C_{\rho,\nu}$ depending on $(\rho,\nu)$.
@@
Then, it is not hard to check the arguments for bounding $\log\mathsf{P}(\mathbf{Y}_{2})$ and $\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})$ directly carry over here, which give the following (corresponding to the earlier (365) and (368)):
$\log\mathsf{P}(\mathbf{Y}_{2})\leq-[1+o(1)+c_{3,\rho,\nu}\delta]ne^{-\nu(1-\frac{\ell}{k})}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho),$ (377)
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=-ne^{-\nu(1-\frac{\ell}{k})}H_{2}(\rho)\big{(}1+c_{4,\rho,\nu}\delta\big{)},$ (378)
@@
which hold with probability at least $1-\exp\big{(}-C^{\prime}_{\rho,\nu}\delta^{2}n\big{)}$ for some $C^{\prime}_{\rho,\nu}>0$ depending on $(\rho,\nu)$. Starting with (354), we can proceed as in (369)–(373) and obtain
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$ $=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$ (379)
$\geq ne^{-\nu(1-\frac{\ell}{k})}\big{[}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\big{]}-nC_{5,\rho,\nu}\delta(1+o(1))$ (380)
$\geq(1-o(1))I_{\ell}^{n}-2nC_{5,\rho,\nu}\delta$ (381)
$=\Big{(}1-o(1)-\frac{2nC_{5,\rho,\nu}\delta}{I_{\ell}^{n}}\Big{)}I_{\ell}^{n},$ (382)
@@
where in (380) we use (377) and (378), and in (381) we apply Lemma 14 below with $C_{5,\rho,\nu}$ being some constant depending on $(\rho,\nu)$. Thus, given any $\delta_{1}>0$, we can take the above $\delta$ as $\delta=\frac{I_{\ell}^{n}\delta_{1}}{4nC_{5,\rho,\nu}}$ to ensure $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\geq(1-o(1)-\frac{\delta_{1}}{2})I_{\ell}^{n}$, which holds with probability at least $1-4\exp\big{(}-\frac{C^{\prime}_{\rho,\nu}(I_{\ell}^{n})^{2}\delta_{1}^{2}}{n}\big{)}$. The proof of Lemma 11 is now complete.
@@
$ne^{-\nu(1-\frac{\ell}{k})}\big{[}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\big{]}\geq(1-o(1))I_{\ell}^{n},$ (383)
where $I_{\ell}^{n}$ satisfies (347).
@@
Since $\frac{\ell}{k}\to 0$, we have $ne^{-\nu(1-\frac{\ell}{k})}=ne^{-\nu}(1+o(1))$. In view of the desired inequality (383) and the asymptotics of $I_{\ell}^{n}$ in (347), some simple algebra reveals that it suffices to prove
$H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\geq(1-o(1))\frac{\nu\ell(1-2\rho)}{k}\log\frac{1-\rho}{\rho}.$ (384)
@@
$e^{-\frac{\nu\ell}{k}}\star\rho$ $=\rho e^{-\frac{\nu\ell}{k}}+(1-\rho)\big{(}1-e^{-\frac{\nu\ell}{k}}\big{)}$ (385)
$=\rho\Big{(}1-(1+o(1))\frac{\nu\ell}{k}\Big{)}+(1-\rho)(1+o(1))\frac{\nu\ell}{k}$ (386)
$=\rho+(1-2\rho+o(1))\frac{\nu\ell}{k}.$ (387)
Then, we only need to apply a Taylor expansion to $H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)$ with $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ to obtain the desired bound (384). ∎
J.2 Proof of Lemma 10 (Mutual Information for the Near-Constant Weight Design)
@@
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}):=\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{Y}}(\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}))$ (388)
@@
$I_{\ell}^{n}=\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{eq}}}}\big{[}I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\big{]}.$ (389)
Recall that $|s_{\mathrm{dif}}|=\ell$ and $|s_{\mathrm{eq}}|=k-\ell$. We first consider $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ for fixed $\mathbf{x}_{s_{\mathrm{eq}}}$, which we expand as (see (18))
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (390)
$=H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-nH_{2}(\rho),$ (391)
where (391) holds because with the full conditioning on $\mathbf{X}_{s}$, the only remaining uncertainty is the i.i.d. Bernoulli($\rho$) noise. It remains to bound $H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}).$ Given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, let $\mathbf{Y}_{1}$ be the collection of outcomes of tests containing at least one item from $s_{\mathrm{eq}}$, let $\mathbf{Y}_{2}$ be the remaining test outcomes, and let the corresponding lengths of $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ be $n_{1}$ and $n_{2}$, respectively. Note also that given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, $\mathbf{Y}_{1}$ only depends on the corresponding $n_{1}$ noise terms, so the vectors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ are conditionally independent. Thus, we get
$H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})+H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (392)
$=n_{1}H_{2}(\rho)+H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}),$ (393)
where (393) is again because under given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, $\mathbf{Y}_{1}$ only depends on the randomness of the $n_{1}$ noise variables. Substituting (393) into (391) yields
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})=H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-n_{2}H_{2}(\rho).$ (394)
@@
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}H(Y^{\prime}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (395)
@@
We use (395) to upper bound the conditional entropy $H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$. Without noise, a given test in $\mathbf{Y}_{2}$ would be negative if and only if it does not contain any item from $s_{\mathrm{dif}}$. Thus, the conditional probability of a single test in $\mathbf{Y}_{2}$ being negative without noise would be $\big{(}1-\frac{1}{n}\big{)}^{\ell\Delta}=\big{(}1-\frac{1}{n}\big{)}^{\frac{\ell\nu n}{k}}=e^{-\alpha\nu(1+o(1))}$. After noise, the probability of a single test being positive becomes $e^{-\alpha\nu(1+o(1))}\star\rho$ (recalling the notation $a\star b=ab+(1-a)(1-b)$). Thus, the conditional entropy of a single test is $H_{2}(e^{-\alpha\nu(1+o(1))}\star\rho)$, and overall (395) yields
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}H_{2}(e^{-\alpha\nu(1+o(1))}\star\rho).$ (396)
Substituting this into (394), we find that
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}(H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho))(1+o(1)).$ (397)
Since (319) in Lemma 7 gives $\mathbbm{E}(n_{2})=n-\mathbbm{E}(n_{1})=n-(1-e^{-\nu(1-\alpha)}+o(1))n=(e^{-\nu(1-\alpha)}+o(1))n$, by further averaging (397) over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (389), we obtain
$I_{\ell}^{n}\leq ne^{-\nu(1-\alpha)}(H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho))(1+o(1)).$ (398)
For the lower bound, we define $V$ to be the number of positive tests in $\mathbf{Y}_{2}$ and first count $(n_{2},V)$ as in first part of the proof of Lemma 11 (Appendix J.1). In particular, for any given $\delta>0$, (357) states that
$n_{2}=\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime}|<\delta$ (399)
with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$; on this event, (363) further gives with probability at least $1-2\exp(-e^{-\nu(1-\alpha)}\delta^{2}n)$ that
$V=ne^{-\nu(1-\alpha)}\big{(}\rho\star e^{-\nu\alpha}+o(1)+c_{\rho,\nu,\alpha}\delta\big{)},\leavevmode\nobreak\ \text{for some }c_{\rho,\nu,\alpha}\leq c_{1}(\rho,\nu,\alpha),$ (400)
@@
$\mathscr{A}_{1}=\big{\{}n_{2}=ne^{-\nu(1-\alpha)}(1+o(1))\big{\}},$ (401)
$\mathscr{A}_{2}=\big{\{}V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))\big{\}},$ (402)

### 23 Oct 2024 v2

@@ I.2 Near-Constant Weight Design
To get started, we note that (J.5) remains valid when $\alpha=0$, which yields that with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$, we have
$n_{2}=\big{(}e^{-\nu(1-\frac{\ell}{k})}+o(1)-\delta^{\prime}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime}|<\delta.$ (J.23)
@@
Next, we note that (J.7) remains valid, which together with (J.23) yields the following with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$:
$M=\big{(}e^{-\nu(1-\frac{\ell}{k})}-e^{-\nu}+o(1)+\delta^{(3)}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{(3)}|<2\delta.$ (J.24)
@@
Moreover, via the randomness of noise, Hoeffding’s inequality still gives (J.9), in which we set $t=n_{2}\delta$, substitute (J.23) and (J.24), and then perform some simple algebra (as in (J.10)–(J.12)) to obtain that
$V=ne^{-\nu(1-\frac{\ell}{k})}\big{(}\rho\star e^{-\frac{\nu\ell}{k}}+o(1)+c_{1,\rho,\nu}\delta\big{)}$ (J.25)
@@
holds with probability at least $1-2\exp(-e^{-\nu}\delta^{2}n)$, where $c_{1,\rho,\nu}$ is a quantity bounded by a constant only depending on $(\rho,\nu)$. By a union bound, the events (J.23), (J.24) and (J.25) collectively hold with probability at least $1-\exp(-C_{\rho,\nu}\delta^{2}n)$ for some $C_{\rho,\nu}$ depending on $(\rho,\nu)$.
@@
Then, it is not hard to check the arguments for bounding $\log\mathsf{P}(\mathbf{Y}_{2})$ and $\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})$ directly carry over here, which give the following (corresponding to the earlier (J.14) and (J.17)):
$\log\mathsf{P}(\mathbf{Y}_{2})\leq-[1+o(1)+c_{3,\rho,\nu}\delta]ne^{-\nu(1-\frac{\ell}{k})}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho),$ (J.26)
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=-ne^{-\nu(1-\frac{\ell}{k})}H_{2}(\rho)\big{(}1+c_{4,\rho,\nu}\delta\big{)},$ (J.27)
@@
which hold with probability at least $1-\exp\big{(}-C^{\prime}_{\rho,\nu}\delta^{2}n\big{)}$ for some $C^{\prime}_{\rho,\nu}>0$ depending on $(\rho,\nu)$. Starting with (J.3), we can proceed as in (J.18)–(J.22) and obtain
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$ $=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$ (J.28)
$\geq ne^{-\nu(1-\frac{\ell}{k})}\big{[}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\big{]}-nC_{5,\rho,\nu}\delta(1+o(1))$ (J.29)
$\geq(1-o(1))I_{\ell}^{n}-2nC_{5,\rho,\nu}\delta$ (J.30)
$=\Big{(}1-o(1)-\frac{2nC_{5,\rho,\nu}\delta}{I_{\ell}^{n}}\Big{)}I_{\ell}^{n},$ (J.31)
@@
where in (J.29) we use (J.26) and (J.27), and in (J.30) we apply Lemma J.1 below with $C_{5,\rho,\nu}$ being some constant depending on $(\rho,\nu)$. Thus, given any $\delta_{1}>0$, we can take the above $\delta$ as $\delta=\frac{I_{\ell}^{n}\delta_{1}}{4nC_{5,\rho,\nu}}$ to ensure $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\geq(1-o(1)-\frac{\delta_{1}}{2})I_{\ell}^{n}$, which holds with probability at least $1-4\exp\big{(}-\frac{C^{\prime}_{\rho,\nu}(I_{\ell}^{n})^{2}\delta_{1}^{2}}{n}\big{)}$. The proof of Lemma H.3 is now complete.
@@
$ne^{-\nu(1-\frac{\ell}{k})}\big{[}H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\big{]}\geq(1-o(1))I_{\ell}^{n},$ (J.32)
where $I_{\ell}^{n}$ satisfies (H.8).
@@
Since $\frac{\ell}{k}\to 0$, we have $ne^{-\nu(1-\frac{\ell}{k})}=ne^{-\nu}(1+o(1))$. In view of the desired inequality (J.32) and the asymptotics of $I_{\ell}^{n}$ in (H.8), some simple algebra reveals that it suffices to prove
$H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)\geq(1-o(1))\frac{\nu\ell(1-2\rho)}{k}\log\frac{1-\rho}{\rho}.$ (J.33)
@@
$e^{-\frac{\nu\ell}{k}}\star\rho$ $=\rho e^{-\frac{\nu\ell}{k}}+(1-\rho)\big{(}1-e^{-\frac{\nu\ell}{k}}\big{)}$ (J.34)
$=\rho\Big{(}1-(1+o(1))\frac{\nu\ell}{k}\Big{)}+(1-\rho)(1+o(1))\frac{\nu\ell}{k}$ (J.35)
$=\rho+(1-2\rho+o(1))\frac{\nu\ell}{k}.$ (J.36)
Then, we only need to apply a Taylor expansion to $H_{2}(e^{-\frac{\nu\ell}{k}}\star\rho)-H_{2}(\rho)$ with $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ to obtain the desired bound (J.33). ∎
J.2 Proof of Lemma H.2 (Mutual Information for the Near-Constant Weight Design)
@@
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}):=\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{Y}}(\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}))$ (J.37)
@@
$I_{\ell}^{n}=\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{eq}}}}\big{[}I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\big{]}.$ (J.38)
Recall that $|s_{\mathrm{dif}}|=\ell$ and $|s_{\mathrm{eq}}|=k-\ell$. We first consider $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ for fixed $\mathbf{x}_{s_{\mathrm{eq}}}$, which we expand as (see (3.1))
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (J.39)
$=H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-nH_{2}(\rho),$ (J.40)
where (J.40) holds because with the full conditioning on $\mathbf{X}_{s}$, the only remaining uncertainty is the i.i.d. Bernoulli($\rho$) noise. It remains to bound $H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}).$ Given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, let $\mathbf{Y}_{1}$ be the collection of outcomes of tests containing at least one item from $s_{\mathrm{eq}}$, let $\mathbf{Y}_{2}$ be the remaining test outcomes, and let the corresponding lengths of $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ be $n_{1}$ and $n_{2}$, respectively. Note also that given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, $\mathbf{Y}_{1}$ only depends on the corresponding $n_{1}$ noise terms, so the vectors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ are conditionally independent. Thus, we get
$H(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})+H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (J.41)
$=n_{1}H_{2}(\rho)+H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}),$ (J.42)
where (J.42) is again because given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$, $\mathbf{Y}_{1}$ only depends on the randomness of the $n_{1}$ noise variables. Substituting (J.42) into (J.40) yields
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})=H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})-n_{2}H_{2}(\rho).$ (J.43)
@@
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}H(Y^{\prime}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (J.44)
@@
We use (J.44) to upper bound the conditional entropy $H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$. Without noise, a given test in $\mathbf{Y}_{2}$ would be negative if and only if it does not contain any item from $s_{\mathrm{dif}}$. Thus, the conditional probability of a single test in $\mathbf{Y}_{2}$ being negative without noise would be $\big{(}1-\frac{1}{n}\big{)}^{\ell\Delta}=\big{(}1-\frac{1}{n}\big{)}^{\frac{\ell\nu n}{k}}=e^{-\alpha\nu(1+o(1))}$. After noise, the probability of a single test being positive becomes $e^{-\alpha\nu(1+o(1))}\star\rho$ (recalling the notation $a\star b=ab+(1-a)(1-b)$). Thus, the conditional entropy of a single test is $H_{2}(e^{-\alpha\nu(1+o(1))}\star\rho)$, and overall (J.44) yields
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}H_{2}(e^{-\alpha\nu(1+o(1))}\star\rho).$ (J.45)
Substituting this into (J.43), we find that
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}(H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho))(1+o(1)).$ (J.46)
Since (G.12) in Lemma G.3 gives $\mathbbm{E}(n_{2})=n-\mathbbm{E}(n_{1})=n-(1-e^{-\nu(1-\alpha)}+o(1))n=(e^{-\nu(1-\alpha)}+o(1))n$, by further averaging (J.46) over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (J.38), we obtain
$I_{\ell}^{n}\leq ne^{-\nu(1-\alpha)}(H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho))(1+o(1)).$ (J.47)
For the lower bound, we define $V$ to be the number of positive tests in $\mathbf{Y}_{2}$ and first count $(n_{2},V)$ as in first part of the proof of Lemma H.3 (Appendix J.1). In particular, for any given $\delta>0$, (J.6) states that
$n_{2}=\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime}|<\delta$ (J.48)
with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$; on this event, (J.12) further gives with probability at least $1-2\exp(-e^{-\nu(1-\alpha)}\delta^{2}n)$ that
$V=ne^{-\nu(1-\alpha)}\big{(}\rho\star e^{-\nu\alpha}+o(1)+c_{\rho,\nu,\alpha}\delta\big{)},\leavevmode\nobreak\ \text{for some }c_{\rho,\nu,\alpha}\leq c_{1}(\rho,\nu,\alpha),$ (J.49)
@@
$\mathscr{A}_{1}=\big{\{}n_{2}=ne^{-\nu(1-\alpha)}(1+o(1))\big{\}},$ (J.50)
$\mathscr{A}_{2}=\big{\{}V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))\big{\}},$ (J.51)